You are asked to prepare a buffer solution with a pH of 3.50. The following solutions, all 0.100 M, are available to you: HCOOH, CH3COOH, H3PO4 , NaCHOO, NaCH3COO, and NaH2PO4

What would be the best combination to make the required buffer solution?


What ratio of [conjugate base]/[acid] would yield the best result for the required buffer solution created in the question above?


1
Expert's answer
2021-07-05T02:34:02-0400

HCOOH and HCOONa


The best ratio is 0.569

PH=pKa +log (HCOONa/HCOOH)

3.5= log (1.8× 10^-4) + log (HCOONa/HCOOH)

Log (HCOONa/ HCOOH)= 3.5 - 3.745

HCOONa/HCOOH =0.569